solve the system of equations by graphing
3x-2y=-6
-3x-y=-12

Answer by mananth(16949)

3x-2y=-6 .............1
-3x-y=-12.............2

equation1
when x= 0 y= 3 ( 0 , 3 )
when y = 0 x= -2 ( -2 , 0 )
Equation 2
when x= 0 y= 12 ( 0 , 12 ) 12
when y = 0 x= 4 ( 4 , 0 )


By graph (2,6) is the solution

m.ananth@hotmail.ca